# Tuning Vision-Language Models with Candidate Labels by Prompt Alignment

### v2

@@ 4.3 The Proposed Framework
Our framework is proposed to further improve the performance of prompt learning with candidate labels. It provides a simple but significant regularization that aligns the mixed prediction of the handcrafted and soft prompt with the current model output using weighted cross-entropy loss. It is shown in Figure 2.
@@
Prompt Alignment Regularization. Let $\mathcal{X}$ be the input space, and $\mathcal{Y}=\{1,2,\ldots,C\}$ be the label space. The $i$-th learnable prompt $\boldsymbol{s}_{i}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},...,\boldsymbol{v}_{M},\boldsymbol{c}_{i}\}$, where $\{\boldsymbol{v}_{m}\}_{m=1}^{M}$ denotes $M$ learnable tokens and $\boldsymbol{c}_{i}$ is the the word embedding for the $i$-th class name. Similarly, the $i$-th manually crafted prompt is denoted as $\boldsymbol{h}_{i}$. To clarify, $f_{i}(\boldsymbol{x};\boldsymbol{s})$ and $g_{i}(\boldsymbol{x};\boldsymbol{h})$ are the softmax outputs of the $i$-th label, as predicted by the learnable prompt and handcrafted prompt separately. When $(\boldsymbol{x},Y)$ is drawn from the partialized dataset, the alignment loss is calculated as:
@@
where $\tilde{p}_{i}$ is mixed linearly with the class posteriors predicted by both prompts:
@@
This regularization term can be adapted to any PLL training objective as:
$\mathcal{L}_{\mathrm{total}}(\boldsymbol{x},Y)=\mathcal{L}_{\mathrm{PLL}}(\boldsymbol{x},Y)+\beta\mathcal{L}_{\mathrm{align}}(\boldsymbol{x},Y).$ (8)
Dynamic Mixing Strategy. In Equation 5, we use a balancing factor $\alpha$ to mix the handcrafted and soft prompt predictions. However, a fixed balancing factor may be sub-optimal since the performance of the soft prompt will surpass the handcrafted prompt as training goes on. With a fixed $\alpha$, the performance of the model will be affected by the low-quality predictions of the handcrafted prompt in the late stages of training. Therefore, a dynamic mixing strategy is adopted to adjust the balancing factor dynamically as:
$\tilde{p}_{i}=\alpha(t)p(y=i\mid\boldsymbol{x};\boldsymbol{s})+(1-\alpha(t))p(y=i\mid\boldsymbol{x};\boldsymbol{h}).$ (9)
Inspired by PLLCR [47], we define $\alpha(t)$ as:
$\alpha(t)=\mathrm{min}\{\frac{t}{T^{\prime}}\lambda,\lambda\}.$ (10)
The model can prioritize learning from the handcrafted prompt at the beginning of training and then focus on the soft prompt subsequently.
@@ 5 Experiments
5.1 Experimental Setting
Datasets.  We adopt $10$ image recognition datasets: ImageNet [5] and Caltech101 [8] for generic object classification; OxfordPets [36], StanfordCars [24], Flowers102 [35], FGVCAircraft [31], and Food101 [1] for fine-grained classification; SUN397 [48] for scene recognition; UCF101 [40] for action classification; DTD [3] for texture classification. In order to evaluate the model’s performance for partial-label learning, as the same in Section 4.1, we define the level of label ambiguity $q$ as the uniform probability of flipping negative labels $\bar{y}_{i}\neq y_{i}$ to false-positive labels inside the candidate label set $Y_{i}$: $q=\mathrm{Pr}(\bar{y}_{i}\in Y_{i}|\bar{y}_{i}\neq y_{i})$. In addition, we use a 16-shot fine-tuning strategy, randomly choosing $16$ images per class from the partialized dataset.
Implementation Details. Our implementation is based on Pytorch [37]. We apply prompt learning on a pre-trained CLIP whose backbone of the image encoder is ResNet-50 [15]. The total number of the learnable prompt tokens is $16$, and the fixed class tokens are at the end of the prompt. Models are trained with a batch size of $32$ and $50$ total epochs for each method and dataset, except for ImageNet, which sets the batch size to $256$. The optimizer is SGD with a cosine decay schedule annealing the learning rate to $0.00001$. The learning rate for prompt learning and our framework is initialized to be $0.002$. Following CoOp[57], the learnable vectors are initialized from a zero-mean Gaussian distribution with a standard deviation equal to 0.02. As for the handcrafted prompts of our framework, we follow the result of prompt engineering of CLIP [38]. For the hyperparameters of our method, we set $\lambda=0.5$, $T^{\prime}=25$, $\beta=1$. Moreover, if a confidence matrix is required in the PLL method, it will be initialized with the model output before training. We conduct all experiments on eight NVIDIA RTX 3090 GPUs and report the average test accuracy and the standard deviation of 4 experiments while keeping the seeds fixed.
5.2 Main Results
PLL Training Objectives. We prove the effectiveness of our framework by incorporating our framework with six state-of-the-art PLL methods on ten datasets. PRODEN [30] progressively identifies the true label and updates the model parameters in a seamless manner. CC [11] assumes that the generation process of candidate labels is uniform and is classifier-consistent. LW [45] weights the loss function by considering the trade-off between losses on candidate labels and the non-candidate labels. PiCO [44] divides the learning process into two components: representation learning by contrastive loss and label disambiguation by prototype and pseudo target updating. PLLCR [47] revisits the idea of consistency regularization and performs supervised learning on non-candidate labels and consistency regularization on candidate labels. CAVL [53] utilizes class activation value (CAV) learning that selects the true label with the maximum CAV. The detailed hyperparameter settings and configurations of the above methods are provided in the supplementary material.
Table 1: Performance comparison of prompt learning and our framework. We evaluate the test accuracy of these two methods with different PLL training objectives at different levels of label ambiguity. Supervised means prompt learning with true labels.
Caltech101 DTD FGVCAircraft Flowers102
Supervised 92.03 62.95 27.27 93.38
@@ 5.2 Main Results
Results on Ten Benchmark Datasets. In Table 1, we compare our framework with prompt learning using six PLL methods when $q=\{0.1,0.3,0.5\}$. The results of ImageNet and SUN397 are presented in the supplemental material. Our framework has shown consistent improvement over most ambiguous conditions, especially when $q=0.5$. For example, in DTD, with PRODEN, our method has gained 1.10%, 1.70% and 10.52%, respectively. Furthermore, in some cases, our framework achieves comparable accuracy with supervised learning, while prompt learning sometimes under-performs zero-shot inference when ambiguity is high. For instance, in OxfordPets with LW at $q=0.5$, prompt learning under-performs zero-shot inference by 5.84%, while our method outperforms supervised learning by 2.21%. In addition, in some datasets, the test accuracy of our framework will increase with more label ambiguity, like in Food101 and OxfordPets. We believe these two fine-grained datasets contain considerable label noise [1, 36], which is supported by the result in CoOp [57], so more label ambiguity actually makes the true label more likely in the candidate label set. Since our framework is more capable of dealing with highly ambiguous candidate labels, more ambiguous candidate labels have exerted a relatively more positive effect on our framework. Notably, there are a few cases where prompt learning with CC and PLLCR has outperformed our method. We conjecture that because these two PLL objectives are considered average-based, whose prediction is not based on the prediction from the last epoch, they suffer less from the error accumulation problem and will not benefit from our framework as effectively as other training objectives.
5.3 Further Analysis
In this part, we conduct extensive experiments to assess the effectiveness and stability of our framework by comparing test accuracy with the PLL training objective of PiCO in 3 representative datasets: DTD, FGVCAircraft and Caltech101.
@@ 5.3 Further Analysis
Influence of Different Handcrafted Prompts.  Because our framework incorporates handcrafted prompts, it is crucial to determine the performance of our framework with differently crafted prompts. In Table 2, we design some prompts with zero-shot performance lower than the default handcrafted prompts [38] and evaluate our framework with these prompts at different levels of label ambiguity. Our framework performs better with better-handcrafted prompts. Better handcrafted prompts have resulted in better test accuracy in our experiments, which indicates that the handcrafted prompts indeed guide the fine-tuning process with candidate labels. Our framework is insensitive to differently crafted prompts. Although the performance of our framework is related to different designs of handcrafted prompts, it is not necessary to design the handcrafted prompt intentionally, as the performance of our framework is insensitive to different handcrafted prompts. Our framework outperforms prompt learning by a large margin even when we simply set the handcrafted prompt “<CLS>.”.
Table 3: Performance comparison with different learnable context lengths when $q=0.3$.
@@
Influence of Different Learnable Context Lengths.  We evaluate our framework when $q=0.3$ while varying the learnable context length of the soft prompt from $1$ to $16$. The result is shown in Table 3. Our framework is superior with different learnable context lengths. It enhances the performance by a large margin with any of the selected context lengths. The longer, the better. Both methods perform worst when the learnable context length is 1 and improve with a longer length. We conjecture that it is because the neural network with fewer parameters is better calibrated [14], reducing the contribution of the model’s zero-shot ability to help disambiguate the candidate labels.
Influence of Different Visual Backbones.  We test the effectiveness of our framework with both ResNet-101 and ViT-B/32 across different label ambiguities. The result is shown in Table 4. Our framework outperforms prompt learning with different visual encoders. Both methods perform better with a better backbone. But our framework is more desirable across different backbones.
@@
Effectiveness of the Dynamic Mixing Strategy.  To evaluate the efficacy of our dynamic and mixing strategy, ablation studies are conducted on the dynamic and mixing strategies. $\alpha(t)$ is set using the same hyperparameters as the main experiment. The result is shown in Table 5. Both dynamic mixing strategies enhance the model’s performance. The model with dynamic technique outperforms all others in 6 and ranked second in 2, out of 9 cases, demonstrating its effectiveness. When only one handcrafted or soft prompt is leveraged, the model’s performance declines and cannot surpass the performance achieved when $\alpha=0.7$. This validates the rationality of the mixing strategy. Lastly, when $\alpha=1$, it outperforms models with only soft prompt guidance or no guidance in most cases, which, to some extent, justifies the effectiveness of explicitly adopting the handcrafted prompt to guide the learning process.
Robustness to Instance-dependent Candidate Labels.  To better simulate the realistic annotation process of candidate labels, we introduce instance-dependent candidate labels and evaluate the performance of our framework under them. Specifically, we use CLIP with the handcrafted prompt of “a photo of a <CLS>.” to make the zero-shot inference of each example. Then, we select the pseudo labels with top {10%, 30%, 50%} confidence as instance-dependent candidate labels. We define the proportion as $q$, and the result is shown in Table 6. Our framework is more robust with instance-dependent candidate labels. Both prompt learning and our framework have experienced a significant performance drop since the instance-dependent candidate labels are more challenging for the model to disambiguate than uniform candidate labels. However, our framework is more robust and can maintain stable performance with high label ambiguity.

### v3

@@ 3.2 Our Framework
This framework is proposed to improve the robustness of prompt learning with candidate labels. It provides a powerful regularization that aligns the dynamically mixed prediction of the handcrafted and learnable prompt with the current model output using weighted cross-entropy loss. It is shown in Figure 2.
Prompt Alignment Regularization.
@@
Let $\mathcal{X}$ be the input space, and $\mathcal{Y}=\{1,2,\ldots,C\}$ be the label space. The $i$-th learnable prompt $\boldsymbol{s}_{i}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},...,\boldsymbol{v}_{M},\boldsymbol{c}_{i}\}$, where $\{\boldsymbol{v}_{m}\}_{m=1}^{M}$ denotes $M$ learnable tokens and $\boldsymbol{c}_{i}$ is the the word embedding for the $i$-th class name. Similarly, the $i$-th manually crafted prompt is denoted as $\boldsymbol{h}_{i}$. To clarify, $f_{i}(\boldsymbol{x};\boldsymbol{s})$ and $g_{i}(\boldsymbol{x};\boldsymbol{h})$ are the softmax outputs of the $i$-th label, as predicted by the learnable prompt and handcrafted prompt separately. When $(\boldsymbol{x},Y)$ is drawn from the partialized dataset, the prompt alignment loss can be calculated as:
@@
where $\tilde{p}_{i}$ is mixed linearly with the class posteriors predicted by both the learnable and handcrafted prompts:
@@
This regularization can be adapted to any prevailing PLL training objective:
$\mathcal{L}_{\mathrm{total}}(\boldsymbol{x},Y)=\mathcal{L}_{\mathrm{PLL}}(\boldsymbol{x},Y)+\beta\mathcal{L}_{\mathrm{align}}(\boldsymbol{x},Y),$ (7)
where $\beta$ is the factor controlling the strength of the alignment loss.
Dynamic Mixing Strategy.
In Equation (4), we use a balancing factor $\alpha$ to mix the handcrafted and learnable prompt predictions. However, a fixed balancing factor may be sub-optimal since the prediction quality of the handcrafted prompt will be influenced by several conditions, such as:
  • Label Ambiguity: In cases of high label ambiguity, the handcrafted prompt tends to provide more reliable predictions, as soft prompts are difficult to learn effectively in such scenarios.
  • Dataset Complexity: For datasets with poor CLIP zero-shot performance, reducing reliance on the handcrafted prompt helps avoid steering the learning process in the wrong direction.
  • Training Epoch: If the soft prompt outperforms the handcrafted prompt as training proceeds, it is better to make more use of the soft prompt, ensuring the most effective predictions are leveraged.
Therefore, to further enhance the robustness of our framework against these conditions, a dynamic mixing strategy is adopted to adjust the balancing factor:
$\tilde{p}_{i}=\alpha(t)p(y=i\mid\boldsymbol{x};\boldsymbol{s})+(1-\alpha(t))p(y=i\mid\boldsymbol{x};\boldsymbol{h}),$ (8)
$\alpha(t)=\mathrm{min}\{\frac{t}{T^{\prime}}\lambda,\lambda\},$ (9)
where $T^{\prime}$ controls how quickly the balancing factor grows over time, adjusting the shift between the two outputs’ contributions and $\lambda$ sets the maximum weight that the output of soft prompts can have, ensuring its influence is capped. Together, they dynamically balance the contributions of the two prompts for robustness.
@@ 4 Experiments
Table 1: Performance comparison of vanilla prompt learning and our framework. We evaluate the test accuracy of these two methods for different PLL training objectives at different label ambiguities. Supervised means prompt learning with true labels. The standard deviation is shown in parentheses.
Caltech101 DTD FGVCAircraft ImageNet
Supervised 92.03 62.95 27.27 60.90
@@
4.1 Experimental Setting
Datasets.  We adopt $8$ image recognition datasets: ImageNet [5] and Caltech101 [7] for generic object classification; OxfordPets [30], StanfordCars [21], FGVCAircraft [27], and Food101 [1] for fine-grained classification; UCF101 [34] for action classification; DTD [3] for texture classification. In order to evaluate the model’s performance for partial-label learning, as the same in Section 2.2, we define the level of label ambiguity $q$ as the uniform probability of flipping negative labels $\bar{y}_{i}\neq y_{i}$ to false-positive labels inside the candidate label set $Y_{i}$: $q=\mathrm{Pr}(\bar{y}_{i}\in Y_{i}|\bar{y}_{i}\neq y_{i})$. In addition, we use a 16-shot fine-tuning strategy, randomly choosing $16$ images per class from the partialized dataset.
Implementation Details. Our implementation is based on Pytorch [31]. We apply prompt learning on a pre-trained CLIP whose backbone of the image encoder is ResNet-50 [13]. The total number of the learnable prompt tokens is $16$, and the fixed class tokens are at the end of the prompt. Models are trained with a batch size of $32$ and $50$ total epochs for each method and dataset, except for ImageNet, which sets the batch size to $256$. The optimizer is SGD with a cosine decay schedule annealing the learning rate to $0.00001$. The learning rate for prompt learning and our framework is initialized to be $0.002$. Following CoOp[49], the learnable vectors are initialized from a zero-mean Gaussian distribution with a standard deviation equal to 0.02. As for the handcrafted prompts of our framework, we follow the result of prompt engineering of CLIP [32]. For example, for the dataset of generic object classification, we adopt “a photo of a <CLS>.” as the handcrafted prompt. For the fine-grained dataset, we adopt the prompts that specify the category, such as “a photo of a <CLS>, a type of food.” for Food101. For the hyperparameters of our method, we set $\lambda=0.5$, $T^{\prime}=25$, $\beta=1$. Moreover, if a confidence matrix is required in the PLL method, it will be initialized with the model output before training. We conduct all experiments on eight NVIDIA RTX 3090 GPUs and report the average test accuracy and the standard deviation of 4 experiments with the seeds fixed.
Training Objectives. We prove the effectiveness of our framework by incorporating six state-of-the-art PLL methods: PRODEN [26], CC [10], LW [38], PiCO [37], PLLCR [40] and CAVL [44].
  • PRODEN [26] progressively identifies the true label and seamlessly updates the model parameters.
  • CC [10] assumes that the generation process of candidate labels is uniform and is provably classifier-consistent.
  • LW [38] weights the loss function by considering the trade-off between losses on candidate and non-candidate labels. We use the loss function version of LWC (LW-Cross entropy) and set the weighting parameter: $\alpha=\beta=1$.
  • PiCO [37] divides the learning process into two components: representation learning by contrastive loss and label disambiguation by prototype and pseudo target updating. We set the queue size that stores key embeddings to be $1024$ and the prototype updating parameter $\gamma=0.9$. In addition, we disable contrastive learning only for one epoch since the classifier already has a satisfying initialization at the start.
  • PLLCR [40] revisits the idea of consistency regularization and performs supervised learning on non-candidate labels and consistency regularization on candidate labels. We set $T^{{}^{\prime}}=50$ in the dynamic balancing strategy: $\gamma(t)=\min\left\{\frac{t}{T^{\prime}}\lambda,\lambda\right\}$. Moreover, We set $T^{{}^{\prime}}$ half of the total epochs, which is the same as the original setting in PLLCR [40].
  • CAVL [44] utilizes class activation value (CAV) learning that selects the true label with the maximum CAV.
4.2 Main Results
In Table 1, we compare our framework with vanilla prompt learning for six PLL training objectives on ten benchmark datasets when $q=\{0.1,0.3,0.5\}$. For supervised learning, we use the same settings as vanilla prompt learning, except the cross entropy is used, while for CLIP zero-shot, we use the same prompt as the handcrafted prompt in our framework. It is shown that the performance of vanilla prompt learning drops significantly with higher label ambiguity and sometimes even under-performs zero-shot inference, verifying its lack of robustness. On the contrary, our framework not only shows robustness with highly ambiguous candidate labels but also improves general performance that matches supervised learning. For instance, in OxfordPets with LW at $q=0.5$, vanilla prompt learning under-performs zero-shot inference by 5.84%, while our framework outperforms supervised learning by 2.21%. Moreover, our framework demonstrates consistent improvement across all levels of label ambiguity, not just in cases of high ambiguity. For example, in the Caltech101 dataset, our framework outperforms vanilla prompt learning in 16 out of 18 cases, regardless of the level of label ambiguity and PLL training objectives. Particularly, the more label ambiguity, the more gain in performance with our framework.
In addition, in some datasets, the test accuracy of our framework will increase with more label ambiguity, like in Food101 and OxfordPets. We believe these two fine-grained datasets contain considerable label noise [1, 30], which is supported by the result in CoOp [49], so more label ambiguity actually makes the true label more likely in the candidate label set. Since our framework can better deal with highly ambiguous candidate labels, more ambiguous candidate labels have exerted a relatively more positive effect on our framework.
There are a few cases where prompt learning with CC and PLLCR has outperformed our method. We conjecture that because these two PLL objectives are considered average-based, whose prediction is not based on the prediction from the last epoch, they suffer less from the error accumulation problem and will not benefit from our framework as effectively as other training objectives.
The main result has verified the efficacy of our framework. By aligning with the handcrafted prompt, the model can make more accurate predictions at the initial stage of the learning process. Moreover, the class posterior predicted by the handcrafted prompt is hardly affected by the level of label ambiguity, making the learning process more robust.
4.3 Ablation Studies
In this part, we conduct experiments to assess the effectiveness of our framework with different handcrafted prompts, learnable context lengths, visual backbones, and balancing strategies in 3 datasets: DTD, FGVCAircraft and Caltech101.
Impact of Different Handcrafted Prompts.  Because our framework incorporates handcrafted prompts, it is crucial to determine the performance of our framework with differently crafted prompts. In Table 2, we design some prompts with zero-shot performance lower than the default handcrafted prompts [32] and evaluate our framework with these prompts at different levels of label ambiguity. Introducing a handcrafted prompt can consistently improve the performance, regardless of the quality of the prompt. Better handcrafted prompts have resulted in better test accuracy in our experiments, which indicates that the handcrafted prompts indeed guide the fine-tuning process with candidate labels. Nonetheless, the performance is improved by a large margin even when we simply set the handcrafted prompt “<CLS>.”.
@@ 4.3 Ablation Studies
Impact of Different Learnable Context Lengths.  We evaluate our framework when $q=0.3$ while varying the learnable context length of the learnable prompt from $1$ to $16$. The result is shown in Table 3. Our framework is superior with different learnable context lengths. It enhances the performance by a large margin with any of the selected context lengths. The longer, the better. Both methods perform worst when the learnable context length is 1 and improve with a longer length. We conjecture that it is because the neural network with fewer parameters is better calibrated [12], reducing the contribution of the model’s zero-shot ability to help disambiguate the candidate labels.
Table 3: Performance comparison with different context lengths when $q=0.3$.
@@
Impact of Different Visual Backbones.  We test the effectiveness of our framework with both ResNet-101 and ViT-B/32 across different label ambiguities. The result is shown in Table 4. Our framework outperforms vanilla prompt learning with different visual encoders. Both methods perform better with a better backbone but our framework is more desirable across different backbones. Moreover, the conclusions from the pilot experiment also apply to VLMs with Vision Transformer as the image encoder.
@@
Effectiveness of the Dynamic Mixing Strategy.  This experiment is conducted to evaluate the efficacy of our dynamic and mixing strategy. $\alpha(t)$ is set using the same hyperparameters as the main experiment. The result is shown in Table 5. The dynamic mixing strategy optimizes the model’s performance. The model with dynamic technique outperforms all others in 6 and ranked second in 2, out of 9 cases, demonstrating its effectiveness. When only one handcrafted or learnable prompt is leveraged, the model’s performance declines and cannot surpass the performance achieved when $\alpha=0.7$. This validates the rationality of the mixing strategy. Lastly, when $\alpha=0$, it outperforms models with only learnable prompt guidance ($\alpha=1$) or no guidance (vanilla prompt learning) in most cases, which, to some extent, justifies the effectiveness of explicitly adopting the handcrafted prompt to guide the learning process.
@@
4.4 Further Analysis
In this part, we demonstrate the effectiveness of our framework with instance-dependent candidate labels or various VLM tuning methods. Moreover, we also highlight its superiority over Unsupervised Prompt Learning [14].
Robustness to Instance-dependent Candidate Labels.  To better simulate the realistic annotation process, we introduce instance-dependent candidate labels and evaluate the performance of our framework under them. Specifically, we use CLIP with the handcrafted prompt of “a photo of a <CLS>.” to make the zero-shot inference of each example. Then, we select the pseudo labels with top {10%, 30%, 50%} confidence as instance-dependent candidate labels. We define the proportion as $q$. The result is shown in Table 6. Our framework is more robust with instance-dependent candidate labels. Both vanilla prompt learning and our framework have experienced a significant performance drop since the instance-dependent candidate labels are more challenging for the model than uniform candidate labels. Nevertheless, our framework is more robust and can maintain stable performance with high label ambiguity.
@@ 4.4 Further Analysis
Performance comparison with UPL on candidate labels.  Unsupervised Prompt Learning [14] (UPL) leverages zero-shot predictions to learn from unsupervised data by optimizing hard pseudo-labels with relatively high confidence for each class. While it shares similarities with our framework, UPL does not incorporate candidate labels and fails to address the error accumulation issue inherent in PLL. As a result, it demonstrates weaker performance compared to our approach as shown in Table 7. These findings further highlight the effectiveness of our framework, particularly when handling candidate labels.
Table 7: Performance comparison on 8 datasets. We set the $q=0.1$ in our dataset and use PRODEN as the surrogate PLL objective. For UPL, we follow the implementation settings in the [14] and set the number of learned samples per class to 16.
       UPL        PRODEN ($q=0.1$)        PRODEN+Ours ($q=0.1$)
       Avg. on 8 datasets        65.89$\pm$ 0.94        69.12$\pm$ 0.83        70.84$\pm$ 0.67
Performance with Different VLM Tuning Methods.  In addition to prompt learning, we integrate our framework with various VLM tuning methods to assess their robustness improvements with candidate labels. Specifically, VPT [17] adjusts visual prompts in the shallow layers, CLIP-Adapter [11] incorporates an adapter module that blends pre-trained features in a residual manner, and Maple [19] simultaneously tunes both visual and textual prompts. The visual architecture is ViT/B16 and the PLL training objective is PRODEN. As shown in Table 8, our framework consistently enhances performance across diverse VLM tuning methods, demonstrating its effectiveness and flexibility.
Table 8: Performance comparison with different VLM tuning methods on Caltech101.
           Method            $q=0.1$            $q=0.3$            $q=0.5$
           Adapter            92.55 $\pm$ 0.15            92.40 $\pm$ 0.24            91.92 $\pm$ 0.43
           Adapter+ours            93.13 $\pm$ 0.12            92.87 $\pm$ 0.05            92.62 $\pm$ 0.15
           VPT            93.52 $\pm$ 0.41            92.03 $\pm$ 0.36            89.02 $\pm$ 0.68
           VPT+ours            93.70 $\pm$ 0.25            92.97 $\pm$ 0.44            92.62 $\pm$ 0.33
           MaPLe            95.00 $\pm$ 0.22            94.33 $\pm$ 0.08            92.77 $\pm$ 0.13
           MaPLe+ours            95.15 $\pm$ 0.09            95.00 $\pm$ 0.10            94.55 $\pm$ 0.09